# Soft decision trees for survival analysis

### v1

@@ 1 Introduction
Survival analysis is an important subfield of Statistics that provides various methods to handle such censored data which are frequently encountered in disciplines such as healthcare, finance, and sociology (Flynn, 2012; Eloranta et al., 2021; Lánczky and Győrffy, 2021; Gepp and Kumar, 2008; Brockett et al., 2008; Zhou et al., 2022; Kramer and Berg, 2003; Plank et al., 2008). The aim of survival models is to estimate for a given input vector $\mathbf{x}$ the relative survival function $S_{\mathbf{x}}(t)$, which represents the probability that the event of interest has not occurred to $\mathbf{x}$ at a certain time $t$. Early works in survival analysis either are based on a nonparametric approach, as in Kaplan-Meier curves (Kaplan and Meier, 1958) where no features are considered (i.e., the specific $\mathbf{x}$ is not considered), or assume some particular parametric form based on linear combination of features, as seen in Cox proportional hazard models (Cox, 1972). Later, several sophisticated statistical models have been proposed (see e.g., the semiparametric approach via splines in Gray (1992); Royston and Parmar (2002, 2011); Luo et al. (2016)), but they often require specific assumptions to be met in order to obtain accurate results.
In addition to traditional statistical methods, various machine learning (ML) models have been developed over the years (Evers and Messow, 2008; Ishwaran et al., 2008; Van Belle et al., 2011; Ranganath et al., 2016; Wang et al., 2019), taking advantage of advances in training and optimization techniques. The adaptation of ML methods to properly handle censored data and time estimation has led to alternatives to statistical models that can capture new non-linear relationships between features and survival times. Several “black-box” ML models have been proposed by adapting the original task to take into account survival data (see e.g., Ripley and Ripley (1998); Che et al. (2018); Ching et al. (2018); Giunchiglia et al. (2018); Katzman et al. (2018); Hu et al. (2021)). Although such black-box models are able to achieve high level of accuracy, the output they produce is often not interpretable, which makes them less practical for use in high-stake domain such as healthcare. Over the years, interpretable ML models have also been proposed and adapted to survival analysis.
@@
Early works on decision trees for classification, regression or survival analysis were based on greedy-like algorithms with often a pruning post-processing step. Due to the remarkable progresses in optimization methods and solvers, growing attention has been recently devoted to globally optimized decision trees, i.e., trees whose parameters are simultaneously tuned and which exhibit local or global optimality guarantees. Examples of Mixed-Integer Linear Optimization (MILO) approaches to design deterministic classification and regression trees can be found in (Bertsimas and Dunn, 2019; Aghaei et al., 2024; Tu and Wu, 2024), of nonlinear optimization approaches to train soft classification and regression trees in (Suárez and Lutsko, 1999; Blanquero et al., 2020, 2022), and of MILO ones to build deterministic survival trees in (Bertsimas et al., 2022; Zhang et al., 2024). See (Costa and Pedreira, 2023) for a survey on decision trees and (Carrizosa et al., 2021) for a review on globally optimized classification and regression trees. In some articles the aim is not only to enhance the predictive performance but also to improve interpretability and to consider fairness issues.
@@
The remainder of the paper is organized as follows. In Section 2 we mention previous work on survival trees. In Section 3, we present the new soft survival tree model, the associated formulation for training, and describe the parametric and spline-based semiparametric models considered in this work for estimating the survival functions within the leaf nodes. We also show how our soft survival tree model can be adapted to take into account group fairness. In Section 4, we extend to soft survival trees the node-based decomposition algorithm proposed in (Consolo et al., 2025) for training soft regression trees. In Section 5, we provide an experimental comparison with three established survival tree models across 15 datasets from the literature. Besides demonstrating how soft trees improve interpretability, we show on a job search dataset how group fairness can be taken into account. Finally, Section 6 contains some concluding remarks and future research directions. Details on the datasets and additional material related to the comparative experiments and the benefits of interpretability are included in the Appendices.
@@ 2 Previous work on survival trees
Since the mid 80s, several greedy approaches have been proposed to adapt decision trees to tackle survival tasks (Gordon and Olshen, 1985; Segal, 1988; Ciampi et al., 1988; Davis and Anderson, 1989; Therneau et al., 1990; LeBlanc and Crowley, 1992; Zhang, 1995; Keleş and Segal, 2002; Molinaro et al., 2004; Jin et al., 2004). These methods use a top-down approach to build tree partitions. At each branch node, a measure of survival dissimilarity (e.g., logrank statistic, survival time variance) identifies the best local split. This process is recursively applied to child nodes without reconsidering previous splits, and hence affecting performance. Pruning heuristics are often employed to prevent overfitting. Widely used greedy methods include the extension to survival analysis of Conditional Inference Trees (Hothorn et al., 2006), CART (Breiman et al., 1984), and the survival trees in (LeBlanc and Crowley, 1993) where the branch node splits maximize the separation between survival functions using the logrank test.
@@
In Bertsimas et al. (2022) the authors extend the local search method introduced in (Dunn, 2018) for univariate and multivariate classification and regression trees to the context of survival analysis. They only consider univariate splitting derived from a proportional hazards model, following the likelihood-based approach proposed by LeBlanc and Crowley (1992). At each leaf node the Nelson-Aalen estimator is used as baseline hazard function, and all individuals within the node are assumed to share the same survival function.
In Huisman et al. (2024), the authors extend the dynamic programming approach proposed in (Demirović et al., 2020) for classification trees to univariate survival trees. As done in (Bertsimas et al., 2022), the splitting rules follow the method of LeBlanc and Crowley (1992), with the Nelson-Aalen estimator used as the hazard function at each leaf node. For the specific case of trees of depth $D=2$, the authors introduce a specialized algorithm to enhance scalability.
In Zhang et al. (2024), a dynamic programming approach is proposed for building optimal univariate survival trees. The loss function employed for the survival task is the so-called Integrated Brier Score (IBS). To limit the exploration of the feasible solution space, bounds on the IBS loss function are introduced. At each leaf node, the Kaplan-Meier estimator is applied.
It is worth pointing out that these three recent survival tree approaches make restrictive assumptions concerning different aspects of the survival tree model and of the training error function. In (Bertsimas et al., 2022; Huisman et al., 2024) the type of univariate splits at the branch nodes are based on proportional hazard, assuming that the ratio between the hazard functions of any two data points remains constant over time. In (Zhang et al., 2024) the survival functions at leaf nodes are estimated via Kaplan-Meier, while in (Bertsimas et al., 2022; Huisman et al., 2024) via Nelson-Aalen estimators. Moreover, in (Zhang et al., 2024) the error function in the training formulation only takes into account the IBS calibration measure.
Finally, an evolutionary algorithm with specific mutation and crossover operators is introduced in Kretowska and Kretowski (2024) for constructing multivariate deterministic survival trees, where Kaplan–Meier estimators are adopted. In the error function a term accounting for the complexity of the tree’s topology is added to the IBS measure.
@@ 3.1 Soft survival tree model
As in (Suárez and Lutsko, 1999; Blanquero et al., 2022) for classification and regression tasks, soft survival trees are defined as maximal binary multivariate trees with a fixed depth $D$ where branch nodes have two children and the leaf nodes are positioned at the same depth.
Let $\tau_{L}$ and $\tau_{B}$ denote the sets of leaf and branch nodes, respectively. At each branch node, a sigmoid function is used to determine the probability of routing a data point to the left or right branch. Specifically, for every input vector $\mathbf{x}_{i}$, with $1\leq i\leq N$, and branch node $n\in\tau_{B}$, the probability of following the left branch is defined as
$p_{in}=F\left(\sum_{j=1}^{p}\omega_{jn}x_{ij}-\omega_{0n}\right),$
where the coefficients $\omega_{jn}\in\mathbb{R}$ and the intercept $\omega_{0n}\in\mathbb{R}$ are the decision variables, and $F$ is the sigmoid function $F(v)=\frac{1}{1+e^{-v}}.$ The probability of taking the right branch is $1-p_{in}$. Due to the soft splitting rule induced by the sigmoid function at each branch node, every input vector $\mathbf{x}_{i}$ is associated with a probability of falling into each leaf node $n\in\tau_{L}$:
$\qquad P_{in}=\prod_{n_{l}\in A_{L(n)}}p_{in_{l}}\;\prod_{n_{r}\in A_{R(n)}}(1-p_{in_{r}}),$
where $A_{L(n)}$ denotes the set of ancestor nodes of leaf node $n$ whose left branches belong to the path from the root to $n$, while $A_{R(n)}$ the set of ancestor nodes for the right branches. Figure 1 shows an example of a soft tree of depth $D=2$.
@@
For any input vector $\mathbf{x}$ and any leaf node $n\in\tau_{L}$, we also have, besides the survival function, the related hazard function and the cumulative hazard function, denoted by $h_{\mathbf{x}}(t;\bm{\beta}_{n})$ and $H_{\mathbf{x}}(t;\bm{\beta}_{n})$, respectively. For the sake of completeness, we briefly recall the relationship between the above functions, and we refer the reader to e.g. (Oakes, 1983) for further details.
@@
[…]
The survival function is then expressed in terms of the cumulative hazard function as:
$S_{\mathbf{x}}(t;\bm{\beta}_{n})=e^{-H_{\mathbf{x}}(t;\bm{\beta}_{n})}.$
@@
For any given values of the parameters $\bm{\omega}_{n}$ with $n\in\tau_{B}$ and $\bm{\beta}_{n}$ with $n\in\tau_{L}$, the soft survival tree provides, for any input vector $\mathbf{x}_{i}\in\mathbb{R}^{p}$, $|\tau_{L}|$ potential survival function predictions (one for each leaf node), along with the associated probabilities $P_{in}$ that $\mathbf{x}_{i}$ is assigned to each corresponding leaf node $n$. In line with (Consolo et al., 2025), for each input $\mathbf{x}$ the actual tree prediction is the survival function $\hat{S}_{\mathbf{x}}(t)=S_{\mathbf{x}}(t;\bm{\beta}_{n_{\mathbf{x}}})$ associated with a single leaf node $n_{\mathbf{x}}$. This leaf node is determined by routing $\mathbf{x}$ from the root along the branches with the highest probability, which correspond to the so-called Highest Branch Probability (HBP) path. Figure 2 illustrates how a tree of depth $D=2$ generates predictions for an input vector $\mathbf{x}$, where $p_{\textbf{x}n}$ represents the probability of x being routed through the left branch of node $n$. The HBP path, highlighted in blue, leads to leaf node $n_{\mathbf{x}}=5$, where the predicted survival function is $\hat{S}_{\mathbf{x}}(t)=S_{\mathbf{x}}(t;\bm{\beta}_{5})$.
Specifically, for any given input vector $\mathbf{x}$ the predicted survival function is defined as:
$\hat{S}_{\mathbf{x}}(t)=S_{\mathbf{x}}(t;\bm{\beta}_{n_{\mathbf{x}}})=\sum_{n\in\tau_{L}}\prod_{\ell\in A_{L(n)}}\mathds{1}_{0.5}(p_{\textbf{x}\ell}(\bm{\omega}_{\ell}))\;\prod_{r\in A_{R(n)}}\mathds{1}_{0.5}(1-p_{\textbf{x}r}(\bm{\omega}_{r}))S_{\mathbf{x}}(t;\bm{\beta}_{n}),$ (1)
where $\mathds{1}_{0.5}(v)=1$ if $v\geq 0.5$ while $\mathds{1}_{0.5}(v)=0$ if $v<0.5$. The Soft Survival Tree model defined above will be referred to as SST.
@@
It is worth pointing out that the deterministic way the prediction is defined for every input vector based on a single leaf node guarantees the beneficial conditional computation property (see e.g., (Bengio et al., 2015)), namely, for any input vector the output only depends on the parameters $\bm{\omega}_{n}$ and $\bm{\beta}_{n}$ for the nodes $n$ contained in the corresponding HBP path. The notable computational and statistical benefits of conditional computation include reducing parameter usage for faster inference and acting as a regularizer to enhance the statistical properties of the model (Breiman, 2001; Hastie et al., 2009; Bengio et al., 2015).
@@ 3.2 Training formulation and advantages of soft survival trees
Before presenting the nonlinear optimization formulation that we propose to train the above-mentioned SSTs, we need to specify the loglikelihood to be used in the estimation of the survival functions associated to the leaf nodes.
Assuming non-informative censoring, that is, assuming that the censoring times of individuals are statistically independent from the occurrence times of the considered event [^1] In clinical settings, this assumption may fail when individuals at higher risk of drop out yield censored survival times, often due to worsening disease status. In extreme cases, dropout may occur shortly before death., the negative log-likelihood associated with each leaf node $n\in\tau_{L}$ and each data point $(\mathbf{x}_{i},t_{i},c_{i})$ is defined as:
$L^{-}_{n}({\mathbf{x}_{i}},t_{i},c_{i};\bm{\beta}_{n})=\begin{cases}-log(h_{\mathbf{x}_{i}}(t_{i};\bm{\beta}_{n}))+H_{\mathbf{x}_{i}}(t_{i};\bm{\beta}_{n})\hskip 14.22636pt\text{if }c_{i}=1,\\ H_{\mathbf{x}_{i}}(t_{i};\bm{\beta}_{n})\hskip 108.12054pt\text{if }c_{i}=0.\end{cases}$
In order to train SSTs, we propose to solve the following unconstrained nonlinear optimization problem:
@@
As to interpretability, SSTs offer two key advantages. First, unlike alternative (deterministic) survival tree models that are constrained to assign a single survival function (e.g., through the Kaplan-Meier estimator) to each leaf node for all input vectors routed to it, the SST approach provides greater flexibility by considering distinct survival functions for each leaf node and input vector. Second, the predictions associated with single leaf nodes can be interpreted as clusters, favoring the analysis of survival functions to extract valuable information about the phenomenon under study. In Section 5.3, we illustrate the valuable additional insights that can be gained using SSTs for three real-world datasets.

### v3

@@ 1 Introduction
Survival analysis is an important subfield of Statistics that provides various methods to handle such censored data which are frequently encountered in disciplines such as healthcare, finance, and sociology (flynn2012survival; eloranta2021cancer; lanczky2021web; gepp2008role; brockett2008survival; zhou2022recurrence; kramer2003survival; plank2008high). The aim of survival models is to estimate for a given input vector $\mathbf{x}$ the relative survival function $S_{\mathbf{x}}(t)$, which represents the probability that the event of interest has not occurred to $\mathbf{x}$ at a certain time $t$. Early works in survival analysis either are based on a nonparametric approach, as in Kaplan-Meier curves (kaplan1958nonparametric) where no features are considered (i.e., the specific $\mathbf{x}$ is not considered), or assume some particular parametric form based on linear combination of features, as seen in Cox proportional hazard models (cox1972regression). Later, several sophisticated statistical models have been proposed (see e.g., the semiparametric approach via splines in gray1992flexible; royston2002flexible; royston2011use; luo2016spline), but they often require specific assumptions to be met in order to obtain accurate results.
In addition to traditional statistical methods, various machine learning (ML) models have been developed over the years (evers2008sparse; ishwaran2008random; van2011support; ranganath2016deep; wang2019machine), taking advantage of advances in training and optimization techniques. The adaptation of ML methods to properly handle censored data and time estimation has led to alternatives to statistical models that can capture new non-linear relationships between features and survival times. Several “black-box” ML models have been proposed by adapting the original task to take into account survival data (see e.g., ripley1998neural; che2018recurrent; ching2018cox; giunchiglia2018rnn; katzman2018deepsurv; hu2021transformer). Although such black-box models are able to achieve high level of accuracy, the output they produce is often not interpretable, which makes them less practical for use in high-stake domain such as healthcare. Over the years, interpretable ML models have also been proposed and adapted to survival analysis.
@@
Early works on decision trees for classification, regression or survival analysis were based on greedy-like algorithms with often a pruning post-processing step. Due to the remarkable progresses in optimization methods and solvers, growing attention has been recently devoted to globally optimized decision trees, i.e., trees whose parameters are simultaneously tuned and which exhibit local or global optimality guarantees. Examples of Mixed-Integer Linear Programming (MILP) approaches to design deterministic classification and regression trees can be found in (bertsimas2019machine; aghaei2020learning; TU2024), of nonlinear optimization approaches to train soft classification and regression trees in (suarez1999globally; blanquero2020sparsity; blanquero2022sparse), and of MILP ones to build deterministic survival trees in (bertsimas2022optimal; zhang2024optimal). Recent works have also focused on methods such as dynamic programming, beam search techniques, and the use of different splitting rules, including those derived from linear programming (e.g., brita2025optimal; kiossou2025generic; lin2024improved; ENGUR2024127354). See (costa2022) for a survey on decision trees and (CARRITOP) for a review on globally optimized classification and regression trees. In some articles the aim is not only to enhance the predictive performance but also to improve interpretability and to consider fairness issues.
@@
The remainder of the paper is organized as follows. In Section 2 we mention previous work on survival trees. In Section 3, we present the new soft survival tree model, the associated formulation for training, and describe the parametric and spline-based semiparametric models considered in this work for estimating the survival functions within the leaf nodes. We also show how our soft survival tree model can be adapted to take into account group fairness. In Section 4, we extend to soft survival trees the node-based decomposition algorithm proposed in (consolo2025) for training soft regression trees. In Section 5, we provide an experimental comparison with three established survival tree models across 15 datasets from the literature. Besides demonstrating how soft trees improve interpretability, we show on a job search dataset how group fairness can be taken into account. Finally, Section 6 contains some concluding remarks and future research directions. Details on the datasets and additional material related to the comparative experiments and the benefits of interpretability are included in the Appendices.
@@ 2 Previous work on survival trees
Since the mid 80s, several greedy approaches have been proposed to adapt decision trees to tackle survival tasks (gordon1985tree; segal1988regression; ciampi1988recpam; davis1989exponential; therneau1990martingale; leblanc1992relative; zhang1995splitting; kelecs2002residual; molinaro2004tree; jin2004alternative). These methods use a top-down approach to build tree partitions. At each branch node, a measure of survival dissimilarity (e.g., logrank statistic, survival time variance) identifies the best local split. This process is recursively applied to child nodes without reconsidering previous splits, and hence affecting performance. Pruning heuristics are often employed to prevent overfitting. Widely used greedy methods include the extension to survival analysis of Conditional Inference Trees (hothorn2006unbiased), CART (breimanclassification), and the survival trees in (leblanc1993survival) where the branch node splits maximize the separation between survival functions using the logrank test.
@@
In bertsimas2022optimal the authors extend the local search method introduced in (dunn2018optimal) for univariate and multivariate classification and regression trees to the context of survival analysis. They only consider univariate splitting derived from a proportional hazards model, following the likelihood-based approach proposed by leblanc1992relative. At each leaf node the Nelson-Aalen estimator is used as baseline hazard function, and all individuals within the node are assumed to share the same survival function.
In huisman2024optimal, the authors extend the dynamic programming approach proposed in (Demirovic2020MurTreeOD) for classification trees to univariate survival trees. As done in (bertsimas2022optimal), the splitting rules follow the method of leblanc1992relative, with the Nelson-Aalen estimator used as the hazard function at each leaf node. For the specific case of trees of depth $D=2$, the authors introduce a specialized algorithm to enhance scalability.
In zhang2024optimal, a dynamic programming approach is proposed for building optimal univariate survival trees. The loss function employed for the survival task is the so-called Integrated Brier Score (IBS). To limit the exploration of the feasible solution space, bounds on the IBS loss function are introduced. At each leaf node, the Kaplan-Meier estimator is applied.
It is worth pointing out that these three recent survival tree approaches make restrictive assumptions concerning different aspects of the survival tree model and of the training error function. In (bertsimas2022optimal; huisman2024optimal) the type of univariate splits at the branch nodes are based on proportional hazard, assuming that the ratio between the hazard functions of any two data points remains constant over time. In (zhang2024optimal) the survival functions at leaf nodes are estimated via Kaplan-Meier, while in (bertsimas2022optimal; huisman2024optimal) via Nelson-Aalen estimators. Therefore, all these models impose the same piecewise-constant survival function on all data points in a leaf node. It is important to note that when only a few data points are contained in the leaf node, this survival function has a small number of large jumps, which may result in an exaggerated drop in survival probability from $t$ to $t+\Delta$ without strong justification. Moreover, in (zhang2024optimal) the error function in the training formulation only takes into account the IBS calibration measure.
Finally, an evolutionary algorithm with specific mutation and crossover operators is introduced in kretowska2024global for constructing multivariate deterministic survival trees, where Kaplan–Meier estimators are adopted. In the error function a term accounting for the complexity of the tree’s topology is added to the IBS measure.
@@ 3.1 Soft survival tree model
As in (suarez1999globally; blanquero2022sparse) for classification and regression tasks, soft survival trees are defined as maximal binary multivariate trees with a fixed depth $D$, where branch nodes have two children and the leaf nodes are positioned at the same depth.
Let $\tau_{L}$ and $\tau_{B}$ denote the sets of leaf and branch nodes, respectively. At each branch node, the probabilities that a data point is routed along the left or right branches are defined as follows.
Definition 1
For each branch node $n\in\tau_{B}$, the probability that an input vector $\mathbf{x}_{i}$, with $1\leq i\leq N$, is routed along the left branch of $n$ is defined as
$p_{in}=F\left(\sum_{j=1}^{p}\omega_{jn}x_{ij}-\omega_{0n}\right)\hskip-3.69899pt,$ (1)
where $F(v)=\frac{1}{1+e^{-v}}$ is a sigmoid function, and the coefficients $\omega_{jn}\in\mathbb{R}$, with $1\leq j\leq p$, and $\omega_{0n}\in\mathbb{R}$ are the decision variables. Obviously, the probability of $\mathbf{x}_{i}$ being routed along the right branch of $n$ is $1-p_{in}$.
Let $A_{L(n)}$ denote the set of ancestor nodes of a leaf node $n$ whose left branches are included in the unique path from the root to $n$, and $A_{R(n)}$ the set of ancestor nodes whose right branches belong to that path. Due to the soft splitting rule (1), the probability that an input vector $\mathbf{x}_{i}$ falls into a given leaf node is defined as follows.
Definition 2
For each leaf node $n\in\tau_{L}$, the probability that an input vector $\mathbf{x}_{i}$, with $1\leq i\leq N$, falls into $n$ is
$P_{in}=\prod_{n_{l}\in A_{L(n)}}p_{in_{l}}\;\prod_{n_{r}\in A_{R(n)}}(1-p_{in_{r}}),$ (2)
namely, it amounts to the product of the (independent) probabilities of $\mathbf{x}_{i}$ being routed along the branches contained in the path from the root to $n$.
An illustrative example of a soft tree of depth $D=2$ is shown in Figure 1.
@@
For any input vector $\mathbf{x}$ and any leaf node $n\in\tau_{L}$, we also have, besides the survival function, the related hazard function and the cumulative hazard function, denoted by $h_{\mathbf{x}}(t;\bm{\beta}_{n})$ and $H_{\mathbf{x}}(t;\bm{\beta}_{n})$, respectively. For the sake of completeness, we briefly recall the relationship between the above functions, and we refer the reader to e.g. (OAKES19833) for further details.
@@
$H_{\mathbf{x}}(t;\bm{\beta}_{n})=\int_{0}^{t}h_{\mathbf{x}}(s;\bm{\beta}_{n})ds.$ (3)
The survival function is then expressed in terms of the cumulative hazard function (3) as:
$S_{\mathbf{x}}(t;\bm{\beta}_{n})=e^{-H_{\mathbf{x}}(t;\bm{\beta}_{n})}.$ (4)
@@
For any given values of the parameters $\bm{\omega}_{n}$ with $n\in\tau_{B}$ and $\bm{\beta}_{n}$ with $n\in\tau_{L}$, the soft survival tree provides, for any input vector $\mathbf{x}_{i}\in\mathbb{R}^{p}$, $|\tau_{L}|$ potential survival function predictions (one for each leaf node), along with the associated probabilities $P_{in}$ that $\mathbf{x}_{i}$ is assigned to each corresponding leaf node $n$. Similarly to what was proposed in (consolo2025) for soft regression trees, for each input $\mathbf{x}$ we define as the actual tree prediction the survival function $\hat{S}_{\mathbf{x}}(t)=S_{\mathbf{x}}(t;\bm{\beta}_{n_{\mathbf{x}}})$ associated with a single leaf node $n_{\mathbf{x}}$. This leaf node is determined by routing $\mathbf{x}$ from the root along the branches with the highest probability, which correspond to the so-called Highest Branch Probability (HBP) path. Figure 2 illustrates how a tree of depth $D=2$ generates the prediction for an input vector $\mathbf{x}$, where $p_{\textbf{x}n}$ represents the probability of x being routed through the left branch of node $n$. The HBP path, highlighted in blue, leads to the leaf node $n_{\mathbf{x}}=5$, where the predicted survival function is $\hat{S}_{\mathbf{x}}(t)=S_{\mathbf{x}}(t;\bm{\beta}_{5})$.
@@
The formal definition of the survival function single leaf node prediction based on the concept of HBP path is as follows.
Definition 3
For a given input vector $\mathbf{x}\in\mathbb{R}^{p}$, the predicted survival function is defined as
$\hat{S}_{\mathbf{x}}(t)=S_{\mathbf{x}}(t;\bm{\beta}_{n_{\mathbf{x}}})=\sum_{n\in\tau_{L}}\prod_{\ell\in A_{L(n)}}\mathds{1}_{0.5}(p_{\mathbf{x}\ell}(\bm{\omega}_{\ell}))\;\prod_{r\in A_{R(n)}}\mathds{1}_{0.5}(1-p_{\mathbf{x}r}(\bm{\omega}_{r}))S_{\mathbf{x}}(t;\bm{\beta}_{n}),$ (5)
where $\mathds{1}_{0.5}(v)=1$ if $v\geq 0.5$ and $\mathds{1}_{0.5}(v)=0$ otherwise.
The Soft Survival Tree model defined above will be referred to as SST.
It is worth pointing out that the deterministic way the prediction is defined for every input vector based on a single leaf node guarantees the beneficial conditional computation property (see e.g., bengio2015conditional), namely, for any input vector the output only depends on the parameters $\bm{\omega}_{n}$ and $\bm{\beta}_{n}$ for the nodes $n$ contained in the corresponding HBP path. The notable computational and statistical benefits of conditional computation include reducing parameter usage for faster inference and acting as a regularizer to enhance the statistical properties of the model (breiman2001random; hastie2009elements; bengio2015conditional).
@@ 3.2 Training formulation and advantages of soft survival trees
Before presenting the nonlinear optimization formulation that we propose to train the above-mentioned SSTs, we need to specify the log-likelihood to be used in the estimation of the survival functions associated to the leaf nodes.
Assuming non-informative censoring, that is, assuming that the censoring times of individuals are statistically independent from the occurrence times of the considered event [^1] In clinical settings, this assumption may fail when individuals at higher risk of drop out yield censored survival times, often due to worsening disease status. In extreme cases, dropout may occur shortly before death. (see, e.g., oquigley2008proportional), the negative log-likelihood associated with each leaf node $n\in\tau_{L}$ and each data point $(\mathbf{x}_{i},t_{i},c_{i})$ is defined as:
$L^{-}_{n}({\mathbf{x}_{i}},t_{i},c_{i};\bm{\beta}_{n})=\begin{cases}-log(h_{\mathbf{x}_{i}}(t_{i};\bm{\beta}_{n}))+H_{\mathbf{x}_{i}}(t_{i};\bm{\beta}_{n})\hskip 14.22636pt\text{if }c_{i}=1,\\ H_{\mathbf{x}_{i}}(t_{i};\bm{\beta}_{n})\hskip 108.12054pt\text{if }c_{i}=0.\end{cases}$ (6)
Based on the definitions provided in (1), (2), and (6) for the leaf and branch nodes, we propose the following unconstrained nonlinear optimization problem to train SSTs:
@@
As to interpretability, SSTs offer two key advantages. First, unlike alternative (deterministic) survival tree models that are constrained to assign a single survival function (e.g., through the Kaplan-Meier estimator) to each leaf node for all input vectors routed to it, the SST approach provides greater flexibility by considering distinct survival functions for each leaf node and input vector. This makes it possible to define a more detailed, data point–specific survival function that relies on the local information associated with the features of each data point. Second, the predictions associated with single leaf nodes can be interpreted as clusters, favoring the analysis of survival functions to extract valuable information about the phenomenon under study. In Section 5.4, we illustrate the valuable additional insights that can be gained using SSTs for three real-world datasets.
Finally, it is worth pointing out that the flexibility of SSTs allows the domain experts to incorporate prior knowledge concerning the application under consideration when selecting an appropriate type of (continuous) survival functions.